OpenStudy (anonymous):

At a certain factory, the daily output is Q(L)= 42000L^(1/3), where L denotes the size of labor force measured in worker-hours. Currently, 1500 worker-hours of labor are used each day. Estimate the effect on output that will be produced if the labor force is (A) cut to 1450 worker-hours (B) increased to 1550 worker-hours

OpenStudy (perl):

was the last answer correct? 27

OpenStudy (anonymous):

yes it was

OpenStudy (perl):

awesome :)

OpenStudy (perl):

i think when it says 'estimate' we should use differentials

OpenStudy (anonymous):

so how do we do that

OpenStudy (perl):

dQ = dQ/dL * dL

OpenStudy (anonymous):

i still don't understand

OpenStudy (perl):

the idea is

OpenStudy (perl):

dy = f'(x) * dx

OpenStudy (perl):

we use dy to approximate the 'actual' change in y

OpenStudy (anonymous):

so we use the derivative of the equation and we dx the equation ?

OpenStudy (perl):

correct

OpenStudy (perl):

I found an expression for dQ/dL, then i multiplied both sides by dL

OpenStudy (anonymous):

and then for b you would just plug in 1550

OpenStudy (aum):

Shouldn't dL be -50 for a) and +50 for b) ?

OpenStudy (perl):

A) dQ = 42000 * 1/3 (1500)^(-2/3) * (-50) B) dQ = 42000 * 1/3 (1500)^(-2/3) * (50)

OpenStudy (anonymous):

why is -50

OpenStudy (perl):

to go from the 'current' 1500 work hours to 1450 , thats -50

OpenStudy (anonymous):

oh

OpenStudy (perl):

dL is the 'increment' in L

OpenStudy (perl):

im not sure what the units are for output, though

OpenStudy (perl):

A) dQ = 42000 * 1/3 (1500)^(-2/3) * (-50) = -5341.999 B) dQ = 42000 * 1/3 (1500)^(-2/3) * (50)=5341.999

OpenStudy (anonymous):

so the answers would be -5320 and 2128

OpenStudy (perl):

I got , rounded to the nearest hundredth or tenth -5342 , 5342

OpenStudy (perl):

Aum saved my butt there

OpenStudy (anonymous):

yes your right

OpenStudy (aum):

np.